what is the probability of rolling an 8 on a number cube?
Answer by checkley71(8403)

assuming you are talking about a pair of die (numbered 1 to 6) where you add up the 2 faces.
then there are 36 possible outcomes & the following combinations add up to 8.
thus we have:
2&6=8
3&5=8
4+4=8
5&3=8
6+2=8
thus there are 5 ways to get an 8
5/36=.139 or 13.9% is the porobability of getting an 8 with the roll of 2 die.